OpenStudy (anonymous):

If x+1/x=2cosθ,then prove that x=cosθ±isinθ

OpenStudy (anonymous):

\[x+\frac{ 1 }{ x }=2\cos \theta\]

OpenStudy (anonymous):

we have to prove that it equals to

OpenStudy (anonymous):

\[x=\cos \theta \pm i \sin \theta\]

hartnn (hartnn):

x^2-2xcos theta +1=0 solve this using formula

OpenStudy (anonymous):

@hartnn i did'nt get it

OpenStudy (shubhamsrg):

note this: we know x+ 1/x >= 2 hence 2cos@ >=2 cos@ >=1 only possibility with real solutions will be cos@ = 1 . for your question, we have on re-arranging , x^2 -2x cos@ + 1=0 x= [2cos@+-sqrt(4cos^2 @ -4) ]/2 = cos@ +- sqrt(-sin^2 @) = required solution..

hartnn (hartnn):

then look ^^

OpenStudy (anonymous):

how did x+1/x=2 ?? i am getting confused ;(

OpenStudy (shubhamsrg):

forget that..on re-arranging, you get a quadratic eqn, just concentrate on that..

OpenStudy (shubhamsrg):

and thats x+ 1/x >= 2 and not =2//but still,,dont think too much about it right now..

OpenStudy (anonymous):

u know i am not used to of these trignometric functions like sin or cos

OpenStudy (anonymous):

thats why i find this question tuff

OpenStudy (anonymous):

where do you get stuck?

OpenStudy (anonymous):

from the start

OpenStudy (anonymous):

i just dont get the logic of question

OpenStudy (anonymous):

first you multiply by x: \[x^2 + 1 = 2\cos \theta x\]

OpenStudy (anonymous):

man i am exploding goku please help

OpenStudy (anonymous):

All it is doing is asking you to solve for x

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

you just have trigonometric terms as the coefficients of your quadratic equation

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

what should i do next after x^2+1=2costhetax

OpenStudy (anonymous):

so you have the equation: \[x^2 -2\cos \theta x +1\] where a=1, b=-2cos(theta), c = 1

OpenStudy (anonymous):

just subtract 2cos(theta)

OpenStudy (anonymous):

so you have the above quadratic, and you can still use the quadratic formula to solve for the roots

OpenStudy (anonymous):

ok wait let me do it

OpenStudy (anonymous):

i am stuck

OpenStudy (anonymous):

I am too learning this Mr moose u good at this well done

OpenStudy (anonymous):

\[\frac{-(-2\cos \theta) \pm \sqrt{(-2\cos \theta)^2 -4}}{2}\]

OpenStudy (anonymous):

can you solve this?

OpenStudy (anonymous):

@Miyuru thank you

OpenStudy (anonymous):

\[x=\frac{ 2\cos \theta \pm \sqrt{4\cos^2 \theta-4} }{2}\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

i am stuck here

OpenStudy (anonymous):

U are well come All the best help @koli123able to solve this i am not good at this hope u the best..

OpenStudy (anonymous):

ok, so lets look at the discriminant: \[4\cos^2\theta-4\] you can factor out a four: \[4(\cos^2\theta -1)\]

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

now, do you remember the Pythagorean identity?

OpenStudy (anonymous):

sin^2 theta

OpenStudy (anonymous):

almost!

OpenStudy (anonymous):

be careful about that

OpenStudy (anonymous):

-sin^2(theta)

OpenStudy (anonymous):

do you think you can figure it out now?

OpenStudy (anonymous):

ok wait

OpenStudy (anonymous):

\[\sqrt{4-\sin^2 \theta }\] stuck here

OpenStudy (anonymous):

\[\frac{2\cos(\theta) \pm \sqrt{-4\sin^2 \theta}}{2}\]

OpenStudy (anonymous):

oh!

OpenStudy (anonymous):

but still i am stuck how will i get its square root

OpenStudy (anonymous):

is it the fact that the square root is imaginary?

OpenStudy (anonymous):

Mr moose is this to be solved using the quadratic formula....can't it be solved using quadratic equation..

OpenStudy (anonymous):

\[2 \sin \theta i\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so cancel out all of the twos and you get...

OpenStudy (anonymous):

got got got it

OpenStudy (anonymous):

Well done @MrMoose u are great/.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

So, in general, when you are solving these types of problems, remember that cos(theta) and sin(theta) and the like can be manipulated like constants

OpenStudy (anonymous):

Also, be sure you know your identities.

OpenStudy (raden):

this is touching :)

OpenStudy (anonymous):

:)